# Efficient Thought Space Exploration through Strategic Intervention

### v1

@@ 1 Introduction
Large language models (LLMs) have made significant advancements in solving complex reasoning tasks (openai_gpt-4_2023; wei_chain--thought_2022; kojima_large_2022; huang_towards_2023), such as mathematical problems (cobbe_training_2021) and code generation (chen_program_2023). The emergence of these reasoning abilities comes from the scaled pre-training (kaplan_scaling_2020; openai2024openaio1card; deepseekai2025deepseekr1incentivizingreasoningcapability) and curated prompting strategies for eliciting reasoning (wei_chain--thought_2022; kojima_large_2022; chen_program_2023; wang_strategic_2024; li2025aisearchparadigm; zheng_progressive-hint_2024).
Recently, researchers have found that leveraging curated decoding strategies to scale LLM inference-time reasoning expansion can also lead to notable performance gains (zhang_what_2025; wang_chain--thought_2024). Typical strategies include extensive sampling, structured rationale topology, and external guidance. For example, self-consistency (wang_self-consistency_2022) and CoT-decoding (wang_chain--thought_2024) sample multiple rationales and select the most consistent answer, Best-of-N (gui_bonbon_2024) uses a score model to select the most reasonable thought from all generations, Tree-of-thoughts (ToT) (yao_tree_2023) and RAP (hao_reasoning_2023) formalize thought exploration as a tree-structure search problem which allows looking ahead and backtracking to make optimal choices, and AdaSwitch (sun_adaswitch_2024) introduces the collaboration with a stronger LLM to overcome the hard part of the reasoning. However, current reasoning exploration methods leave three unsolved issues: (1) Low token efficiency: pure-sampling methods cannot reuse correct prefixes while tree-structured methods generate numerous intermediate branches but only a few of them can lead to the final answer. Furthermore, when dealing with complex problems, these methods require sampling prohibitively many trajectories to find a correct path. (2) Local exploration guidance: heuristic-based search algorithm easily converges to a “good” path and fails to explore the whole thought space. Although MCTS uses UCT to balance exploration and exploitation, it does not fully exploit token probabilities, resulting in an underestimation of truly valuable branching points. (3) High external dependence: to verify generated steps or overcome overly hard reasoning steps, a common practice is to apply an ad-hoc trained verifier or considerable usage of a stronger LLM reasoner, which will introduce considerable extra cost.
@@
Building upon the empirical observation that strategic intervention at sparse critical tokens can effectively redirect reasoning trajectories at minimal cost, we accordingly propose a novel Hint-Practice Reasoning (HPR) framework (illustrated in Figure 1), which implements this insight via two distinct roles: (1) a hinter (powerful but costly model) provides targeted hints at sparse critical tokens; (2) a practitioner (efficient smaller model) executes the majority of reasoning steps. Different from existing multi-agent studies (sun_adaswitch_2024; gui_bonbon_2024), HPR hinter selectively corrects the critical mistakes of the practitioner and additionally guides exploration, which reduces the large model’s generation burden. The left challenge lies in identifying where to apply interventions. Intuitively, the practitioner should think twice when it significantly deviates from hinter’s prediction. Furthermore, we expect the practitioner to start exploring from a promising and uncertain (from the global view) partial trajectory to reuse computation and maximize exploration.
@@ 2.1 Inference-time LLM Reasoning
Recent studies have found increasing inference-time computations can improve the output quality of LLMs, particularly for complex tasks that needs long-chain reasoning (wu_inference_2024; brown_large_2024; fu_efficiently_2024; chu_navigate_2024). Chain-of-thought prompting (wei_chain--thought_2022) is the first to demonstrate intermediate rationales can boost the reasoning ability of LLMs. Zero-shot CoT (kojima_large_2022) employs a reasoning instruction to activate the rationale generation and avoid the dependence on few-shot demonstrations. Subsequent research focuses on further improving the quality of the rationale. UAG (yin_reasoning_2024) identifies the risky steps of the reasoning path by the uncertainty and adjust demonstrations to eliminate the risk. zhou_least--most_2023; press_measuring_2023 prompt the LLM to decompose the query into easier sub-parts. madaan_self-refine_2023; manakul_selfcheckgpt_2023 self-correct the reasoning mistakes by tasks-specific prompting. kumar2024traininglanguagemodelsselfcorrect internalized the self-correction ability by reinforcement learning. cobbe_training_2021; wang_math-shepherd_2024; li_making_2023 propose to train a verifier to select the optimal rationale. Our approach exploit the potentials of autoregressive decoding, which is task-agnostic and does not require curated demonstration, prompts, nor ad-hoc trained models.
@@ 2.2 Thought Space Exploration
The autoregressive decoding nature of LLM poses challenges in finding the globally optimal reasoning path. To address this, researchers utilized the probabilistic modeling nature of LLMs to explore diverse thoughts. Self-consistency (wang_self-consistency_2022) adopts a sample-and-marginalize decoding strategy, finding the most consistent answer among multiple sampled paths. CoT-decoding (wang_chain--thought_2024) create diverse paths by enumerating the top-K candidates of the starting token and select the most plausible path by comparing answer’s confidence score. Further advancing this line of research, Tree-of-thoughts and Graph-of-Thoughts (yao_tree_2023; long_large_2023; besta_graph_2024; mo_tree_2024; zhang_thought_2024) generalizes the topology of rationale into the tree or graph structures by allowing LLM to try multiple reasoning branches at each step. RAP (hao_reasoning_2023) regard reasoning as planning and introduce Monte Carlo Tree Search (MCTS) to organize the reasoning exploration space, which iteratively expand new paths under the guide of reward models. AdaSwitch (sun_adaswitch_2024) formulates rationale generation as a multi-agent task. When the local model finds an error in the reasoning process, it asks for a more powerful cloud model to correct the mistake. Our approach follows tree-structure thought exploration paradigm. We achieve more efficient reasoning by delegating part of generation to a stronger reasoner and further exploiting LLM’s probabilistic information to organize the tree growth.
@@ 3 Sparse Disparity in LLM Reasoning
To give an intuitive illustration, we conduct an analytic experiment on MATH (hendrycks_measuring_2021) dataset that measures the disparity between a small LLM $\phi$ (Qwen2.5-3B-Instruct) and a high-performance oracle $\theta$ (Qwen2.5-32B-Instruct). Specifically, we feed model $\phi$ with an output from model $\theta$ to make predictions at each position, collect the inconsistent ones, and continue generation from the inconsistent positions using $\phi$ to see whether it leads to an incorrect final answer (deviation).
@@ 4.1 Targeted Reasoning Tree Expansion
HPR aims to allocate test-time computation to the reasoning state where it is truly needed. We adopt a tree-structured search to progressively explore the thought space and reuse high-quality rationale prefixes, avoiding redundant computation. To aggregate reasoning paths, we apply weighted voting, a proven method for test-time scaling (zhang_what_2025; wang_self-consistency_2022).
However, prior tree-based methods (yao_tree_2023; mo_tree_2024; hao_reasoning_2023) often generate many incomplete paths, which are excluded from final voting, leading to inefficiency. To address this, we ensure every expansion yields a complete rationale by letting the practitioner finish each branch.
@@ 5.1 Setup
We evaluate our approach on arithmetic reasoning and commonsense reasoning tasks. Arithmetic reasoning includes GSM8K (cobbe_training_2021), AQUA-RAT (chen_program_2023), and MATH (hendrycks_measuring_2021). These datasets are among the more challenging ones used in prior studies (sun_adaswitch_2024; ma_non-myopic_2025). For commonsense reasoning, we select CSQA (talmor_commonsenseqa_2019) and StrategyQA (geva_did_2021), both of which require implicit reasoning.
@@
We use models from Qwen2.5 series (qwen2.5): 3B/7B for practitioner and 14B/32B for hinter. To account for differences in task complexity, we set the HPR hint length to 32 tokens for arithmetic reasoning and 16 tokens for commonsense reasoning. We use vLLM (kwon_efficient_2023) library to provide efficient inference serving. For all experiments, we use “Let’s think step by step” prompting (kojima_large_2022) to perform zero-shot CoT. We set the temperature to 0.7 for all non-greedy generations and average results across three different seeds.
@@
We compare HPR with single-trajectory CoT, sampling-based method Self-Consistency (SC) (wang_self-consistency_2022), tree-structure search methods ToT (yao_tree_2023), TouT (mo_tree_2024), and MCTS (hao_reasoning_2023), as well as externally guided baselines Best-of-N (BoN) (song_good_2024) and AdaSwitch (sun_adaswitch_2024). Besides, we report a pure hinter baseline to show the performance upper bound of externally guided methods. We use the hinter as the reward model both for BoN and MCTS. For MCTS, we follow Marco-o1 (zhao_marco-o1_2024) to use the average confidence as the reward score, and ensemble all paths via majority voting. ToT, TouT, and MCTS are allowed to sample 5 different actions at each step in our experiments. The hyperparameters for AdaSwitch are tuned to ensure a comparable token budget with HPR.
@@
We extract the final answer from the LLM’s output and calculate exact match accuracy. To distinguish the contributions of the practitioner and the hinter, we recorded the average number of tokens involved by each during inference (averaged across samples). We calculate the FLOPs (floating point operations) to measure the cost of different approaches. Following kaplan_scaling_2020, the FLOPs of decoding a sequence of length $L$ approximately equal $2NL$, where $N$ is the number of non-embedding parameters of the LLM. To evaluate the reasoning-expansion efficiency (REE), we devise an integrated measure by computing the normalized gain-cost ratio:
@@ 7 Future Work
Method Number of Paths
CoT-SC 1,3,4,5,7,9,11,13,15
TouT 1,2,3,4,5
MCTS 1,2,3,4,5
AdaSwitch 1,2,3,4,5
HPR 1,2,3,4,5,7 9
Table 4: Number of reasoning paths for each method in the efficiency analysis.

### v2

@@ 1 Introduction
Large language models (LLMs) have made significant advancements in solving complex reasoning tasks (OpenAI 2023; Wei et al. 2022; Kojima et al. 2022; Huang and Chang 2023), such as mathematical problems (Cobbe et al. 2021) and code generation (Chen et al. 2023). The emergence of these reasoning abilities comes from the scaled pre-training (Kaplan et al. 2020; OpenAI 2024; DeepSeek-AI 2025) and curated prompting strategies for eliciting reasoning (Wei et al. 2022; Kojima et al. 2022; Chen et al. 2023; Wang et al. 2024b; Li et al. 2025; Zheng et al. 2024).
Recently, researchers have found that leveraging curated decoding strategies to scale LLM inference-time reasoning expansion can also lead to notable performance gains (Zhang et al. 2025; Wang and Zhou 2024). Typical strategies include extensive sampling, structured rationale topology, and external guidance. For example, self-consistency (Wang et al. 2023) and CoT-decoding (Wang and Zhou 2024) sample multiple rationales and select the most consistent answer, Best-of-N (Gui et al. 2024) uses a score model to select the most reasonable thought from all generations, Tree-of-thoughts (ToT) (Yao et al. 2023) and RAP (Hao et al. 2023) formalize thought exploration as a tree-structure search problem which allows looking ahead and backtracking to make optimal choices, and AdaSwitch (Sun et al. 2024) introduces the collaboration with a stronger LLM to overcome the hard part of the reasoning. However, current reasoning exploration methods leave three unsolved issues: (1) Low token efficiency: pure-sampling methods cannot reuse correct prefixes while tree-structured methods generate numerous intermediate branches but only a few of them can lead to the final answer. Furthermore, when dealing with complex problems, these methods require sampling prohibitively many trajectories to find a correct path. (2) Local exploration guidance: heuristic-based search algorithm easily converges to a “good” path and fails to explore the whole thought space. Although MCTS uses UCT to balance exploration and exploitation, it does not fully exploit token probabilities, resulting in an underestimation of truly valuable branching points. (3) High external dependence: to verify generated steps or overcome overly hard reasoning steps, a common practice is to apply an ad-hoc trained verifier or considerable usage of a stronger LLM reasoner, which will introduce considerable extra cost.
@@
Building upon the empirical observation that strategic intervention at sparse critical tokens can effectively redirect reasoning trajectories at minimal cost, we accordingly propose a novel Hint-Practice Reasoning (HPR) framework (illustrated in Figure 1), which implements this insight via two distinct roles: (1) a hinter (powerful but costly model) provides targeted hints at sparse critical tokens; (2) a practitioner (efficient smaller model) executes the majority of reasoning steps. Different from existing multi-agent studies (Sun et al. 2024; Gui et al. 2024), HPR hinter selectively corrects the critical mistakes of the practitioner and additionally guides exploration, which reduces the large model’s generation burden. The left challenge lies in identifying where to apply interventions. Intuitively, the practitioner should think twice when it significantly deviates from hinter’s prediction. Furthermore, we expect the practitioner to start exploring from a promising and uncertain (from the global view) partial trajectory to reuse computation and maximize exploration.
@@ 2.1 Inference-time LLM Reasoning
Recent studies have found increasing inference-time computations can improve the output quality of LLMs, particularly for complex tasks that needs long-chain reasoning (Wu et al. 2024; Brown et al. 2024; Fu et al. 2024; Chu et al. 2024). Chain-of-thought prompting (Wei et al. 2022) is the first to demonstrate intermediate rationales can boost the reasoning ability of LLMs. Zero-shot CoT (Kojima et al. 2022) employs a reasoning instruction to activate the rationale generation and avoid the dependence on few-shot demonstrations. Subsequent research focuses on further improving the quality of the rationale. UAG (Yin et al. 2024) identifies the risky steps of the reasoning path by the uncertainty and adjust demonstrations to eliminate the risk. Zhou et al. (2023); Press et al. (2023) prompt the LLM to decompose the query into easier sub-parts. Madaan et al. (2023); Manakul et al. (2023) self-correct the reasoning mistakes by tasks-specific prompting. Kumar et al. (2024) internalized the self-correction ability by reinforcement learning. Cobbe et al. (2021); Wang et al. (2024a); Li et al. (2023) propose to train a verifier to select the optimal rationale. Our approach exploit the potentials of autoregressive decoding, which is task-agnostic and does not require curated demonstration, prompts, nor ad-hoc trained models.
@@ 2.2 Thought Space Exploration
The autoregressive decoding nature of LLM poses challenges in finding the globally optimal reasoning path. To address this, researchers utilized the probabilistic modeling nature of LLMs to explore diverse thoughts. Self-consistency (Wang et al. 2023) adopts a sample-and-marginalize decoding strategy, finding the most consistent answer among multiple sampled paths. CoT-decoding (Wang and Zhou 2024) create diverse paths by enumerating the top-K candidates of the starting token and select the most plausible path by comparing answer’s confidence score. Further advancing this line of research, Tree-of-thoughts and Graph-of-Thoughts (Yao et al. 2023; Long 2023; Besta et al. 2024; Mo and Xin 2024; Zhang et al. 2024) generalizes the topology of rationale into the tree or graph structures by allowing LLM to try multiple reasoning branches at each step. RAP (Hao et al. 2023) regard reasoning as planning and introduce Monte Carlo Tree Search (MCTS) to organize the reasoning exploration space, which iteratively expand new paths under the guide of reward models. AdaSwitch (Sun et al. 2024) formulates rationale generation as a multi-agent task. When the local model finds an error in the reasoning process, it asks for a more powerful cloud model to correct the mistake. Our approach follows tree-structure thought exploration paradigm. We achieve more efficient reasoning by delegating part of generation to a stronger reasoner and further exploiting LLM’s probabilistic information to organize the tree growth.
@@ 3 Sparse Disparity in LLM Reasoning
To give an intuitive illustration, we conduct an analytic experiment on MATH (Hendrycks et al. 2021) dataset that measures the disparity between a small LLM $\phi$ (Qwen2.5-3B-Instruct) and a high-performance oracle $\theta$ (Qwen2.5-32B-Instruct). Specifically, we feed model $\phi$ with an output from model $\theta$ to make predictions at each position, collect the inconsistent ones, and continue generation from the inconsistent positions using $\phi$ to see whether it leads to an incorrect final answer (deviation).
@@ 4.1 Targeted Reasoning Tree Expansion
HPR aims to allocate test-time computation to the reasoning state where it is truly needed. We adopt a tree-structured search to progressively explore the thought space and reuse high-quality rationale prefixes, avoiding redundant computation. To aggregate reasoning paths, we apply weighted voting, a proven method for test-time scaling (Zhang et al. 2025; Wang et al. 2023).
However, prior tree-based methods (Yao et al. 2023; Mo and Xin 2024; Hao et al. 2023) often generate many incomplete paths, which are excluded from final voting, leading to inefficiency. To address this, we ensure every expansion yields a complete rationale by letting the practitioner finish each branch.
@@ 5.1 Setup
We evaluate our approach on arithmetic reasoning and commonsense reasoning tasks. Arithmetic reasoning includes GSM8K (Cobbe et al. 2021), AQUA-RAT (Chen et al. 2023), and MATH (Hendrycks et al. 2021). These datasets are among the more challenging ones used in prior studies (Sun et al. 2024; Ma et al. 2025). For commonsense reasoning, we select CSQA (Talmor et al. 2019) and StrategyQA (Geva et al. 2021), both of which require implicit reasoning.
@@
We use models from Qwen2.5 series (Yang et al. 2024): 3B/7B for practitioner and 14B/32B for hinter. To account for differences in task complexity, we set the HPR hint length to 32 tokens for arithmetic reasoning and 16 tokens for commonsense reasoning. We use vLLM (Kwon et al. 2023) library to provide efficient inference serving. For all experiments, we use “Let’s think step by step” prompting (Kojima et al. 2022) to perform zero-shot CoT. We set the temperature to 0.7 for all non-greedy generations and average results across three different seeds.
@@
We compare HPR with single-trajectory CoT, sampling-based method Self-Consistency (SC) (Wang et al. 2023), tree-structure search methods ToT (Yao et al. 2023), TouT (Mo and Xin 2024), and MCTS (Hao et al. 2023), as well as externally guided baselines Best-of-N (BoN) (Song et al. 2024) and AdaSwitch (Sun et al. 2024). Besides, we report a pure hinter baseline to show the performance upper bound of externally guided methods. We use the hinter as the reward model both for BoN and MCTS. For MCTS, we follow Marco-o1 (Zhao et al. 2024) to use the average confidence as the reward score, and ensemble all paths via majority voting. ToT, TouT, and MCTS are allowed to sample 5 different actions at each step in our experiments. The hyperparameters for AdaSwitch are tuned to ensure a comparable token budget with HPR.
@@
We extract the final answer from the LLM’s output and calculate exact match accuracy. To distinguish the contributions of the practitioner and the hinter, we recorded the average number of tokens involved by each during inference (averaged across samples). We calculate the FLOPs (floating point operations) to measure the cost of different approaches. Following Kaplan et al. (2020), the FLOPs of decoding a sequence of length $L$ approximately equal $2NL$, where $N$ is the number of non-embedding parameters of the LLM. To evaluate the reasoning-expansion efficiency (REE), we devise an integrated measure by computing the normalized gain-cost ratio: